# WikiGap: Promoting Epistemic Equity by Surfacing Knowledge Gaps Between English Wikipedia and other Language Editions

### v3

@@ 7. Conclusion
WikiGap contends with longstanding structural inequalities in Wikipedia by enabling readers to access cross-lingual facts directly within the English interface. Through a novel combination of fact-centric text comparison (InfoGap; Samir et al., 2024) based on LLMs (Min et al., 2023) and user-centered design, the system unobtrusively displays knowledge missing due to linguistic and editorial asymmetries. Our evaluation shows that WikiGap not only improves usability and fact-finding efficiency, but also fosters greater awareness of cultural variation and the limitations of assuming English Wikipedia as a default knowledge base. By surfacing multilingual content in context, WikiGap connects readers with perspectives on topics that are situated outside of the Anglosphere, connections that were previously occluded by the hegemonic dominance of English Wikipedia (Kumar, 2021, Chapter 3). This work demonstrates how augmentative tools can shift reader practices and challenge epistemic assumptions, paving the way for more inclusive, transparent, and culturally grounded knowledge infrastructures.
@@ Appendix C Average Completion Time and Accuracy across Different Topics
Figure 8. Side-by-side bar charts comparing average completion time and accuracy across different topics, with and without the WikiGap extension. Blue represent the control condition (no WikiGap), and orange represent the treatment condition (with WikiGap).

### v4

@@ 7. Conclusion
WikiGap contends with longstanding structural inequalities in Wikipedia by enabling readers to access cross-lingual facts directly within the English interface. Through a novel combination of fact-centric text comparison (InfoGap; Samir et al., 2024) based on LLMs (Min et al., 2023) and user-centered design, the system unobtrusively displays knowledge missing due to linguistic and editorial asymmetries. Our evaluation shows that WikiGap not only improves usability and fact-finding efficiency, but also fosters greater awareness of cultural variation and the limitations of assuming English Wikipedia as a default knowledge base. By surfacing multilingual content in context, WikiGap connects readers with perspectives on topics that are situated outside of the Anglosphere, connections that were previously occluded by the hegemonic dominance of English Wikipedia (Kumar, 2021, Chapter 3). This work demonstrates how augmentative tools can shift reader practices and challenge epistemic assumptions, paving the way for more inclusive, transparent, and culturally grounded knowledge infrastructures. This paper examines how interface design can challenge the English-as-superset assumption by making multilingual knowledge differences visible to readers during everyday Wikipedia use.
@@ Appendix C Average Completion Time and Accuracy across Different Topics
Figure 8. Side-by-side bar charts comparing average completion time and accuracy across different topics, with and without the WikiGap extension. Blue represents the control condition (no WikiGap), and orange represent the treatment condition (with WikiGap).